# Diversifying Query: Region-Guided Transformer for Temporal Sentence Grounding

### v1

@@ 4.1 Datasets, Metrics, and Implementation Details
M-DETR [11]NeurIPS’21 52.89 33.02 54.82 29.40 30.73 53.94 34.84 - - 32.20
UMT [19]CVPR’22 56.23 41.18 53.83 37.01 36.12 60.26 44.26 - - 38.59
QD-DETR [24]CVPR’23 62.40 44.98 62.52 39.88 39.86 62.68 46.66 62.23 41.82 41.22
UniVTG [14]ICCV’23 58.86 40.86 57.60 35.59 35.47 59.74 - - - 36.13
EaTR [9]ICCV’23 - - - - - 61.36 45.79 61.86 41.91 41.74
MomentDiff [13]NeurIPS’23 57.42 39.66 54.02 35.73 35.95 - - - - -
TR-DETR [31]AAAI’24 64.66 48.96 63.98 43.73 42.62 67.10 51.48 66.27 46.42 45.09
TaskWeave [38]CVPR’24 - - - - - 64.26 50.06 65.39 46.47 45.38
UVCOM [36]CVPR’24 63.55 47.47 63.37 42.67 43.18 65.10 51.81 - - 45.79
CG-DETR [23]Arxiv‘24 65.43 48.38 64.51 42.77 42.86 67.35 52.06 65.57 45.73 44.93
@@
Method TACoS Charades-STA
R0.3 R0.5 R0.7 mIoU R0.3 R0.5 R0.7 mIoU
2D-TAN [41] 40.01 27.99 12.92 27.22 58.76 46.02 27.50 41.25
M-DETR [11] 37.97 24.67 11.97 25.49 65.83 52.07 30.59 45.54
MomentDiff [13] 44.78 33.68 - - - 55.57 32.42 -
QD-DETR [24] - - - - - 57.31 32.55 -
UniVTG [14] 51.44 34.97 17.35 33.60 70.81 58.01 35.65 50.10
CG-DETR [23] 52.23 39.61 22.23 36.48 70.43 58.44 36.34 50.13
RGTR (Ours) 53.04 40.31 24.32 37.44 72.04 57.93 35.16 50.32
Table 2: Performances on TACoS and Charades-STA. Video features are extracted using Slowfast and CLIP.
@@
Method R0.5 R0.7
2D-TAN [41] 40.94 22.85
FVMR [7] 42.36 24.14
SSRN [43] 46.72 27.98
UMT [19] 48.31 29.25
MomentDiff [13] 51.94 28.25
QD-DETR [24] 52.77 31.13
TR-DETR [31] 53.47 30.81
CG-DETR [23] 55.22 34.19
RGTR (Ours) 55.48 34.33
Table 3: Results on Charades-STA with VGG backbone.
Implementation Details.  Following previous methods [24, 13, 23], for all three datasets, we use SlowFast [5] and CLIP [26] to extract visual features and the text encoder in CLIP to extract textual features. In Charades-STA, we also extract visual features with VGG [29] and use Glove [25] for textual features. The cross-modal alignment encoder and region-guided decoder consist of three layers of transformer blocks. In the encoder, we also use a local regular loss following [31]. We set the embedding dimension $D$ to 256 and the number of attention heads to 8. The number of anchor pairs $\mathcal{K}$ is set to 20 for QVHighlights, 10 for Charades-STA and TACoS. The NMS threshold is set to 0.8. The balancing parameters are set as: $\lambda_{\text{align}}=0.3$, $\lambda_{\text{iou}}=1$, and $\lambda_{\text{sal}}$ is set as 1 for QVHighlights, 4 for Charades-STA and TACoS. We train all the models with batch size 32 for 200 epochs using the AdamW optimizer [20] with weight decay 1e-4 for all three datasets. The learning rate is set to 1e-4.
@@ 4.2 Performance Comparison
As shown in Tab. 1, we compare RGTR to previous methods on QVHighlights. For a fair comparison, we report numbers for both the test and validation splits. Our RGTR achieves new state-of-the-art performance on all metrics. Specifically, RGTR outperforms the latest methods like CG-DETR, achieving 67.12% at mAP@0.5 and 45.77% at mAP@0.75 on the test dataset. Particularly, the average mAP score of 45.53% on the test dataset marks a significant improvement over UVCOM by 2.35%. On the validation dataset, RGTR also maintains its lead. The notable performance advantages of RGTR demonstrate the effectiveness of anchor pairs with explicit regional priors.
Tab. 3 and Tab 3 present comparisons on TACoS and Charades-STA. Our RGTR achieves the best performance on TACoS. On Charades-STA, RGTR also maintains its competitiveness regardless of whether 2D features (VGG) or multimodal pre-trained features (SF+C) are used. However, we observe that while our results are notably superior on QVHighlights, the margin is slightly reduced on TACoS and Charades-STA. We attribute this to the biased distribution on Charades-STA and TACoS compared to QVHighlights, resulting in less diversity learned by anchor pairs. We also provide results of RGTR on the anti-biased Charades-STA in the supplemental material.
Table 4: Ablation study on the components of RGTR on QVHighlights val split.
Explicit Anchor Initialization Region-Guided Attention Module IoU-Aware Scoring Head R1 mAP
  @0.5   @0.7   @0.5   @0.75   Avg.
65.35 48.97 64.58 43.05 43.12
64.65 50.58 64.98 45.50 44.82
66.19 49.61 65.06 44.43 44.03
65.55 51.29 65.96 46.08 45.36
66.13 51.68 66.31 47.55 46.51
67.68 52.90 67.38 48.00 46.95
@@
Distribution R0.5 R0.7 mAPavg
random 66.19 49.61 44.03
uniform grids 67.10 50.97 44.93
k-means 67.68 52.90 46.95
Table 5: Ablation study on distribution of anchor initialization.
@@
Loss type R0.5 R0.7 mAPavg
Huber Loss 68.00 51.94 46.68
L1 Loss 66.65 51.89 46.73
L2 Loss 67.68 52.90 46.95
Table 6: Ablation study on the IoU loss type.
[…]
Scoring R0.5 R0.7 mAPavg
IoU superv. 67.87 52.84 46.54
cls $+$ IoU 67.23 52.39 46.92
cls $\times$ IoU 67.68 52.90 46.95
Table 7: Ablation study on various scoring methods.
[…]
{adjustwidth}
-0.15cm0cm Refer to caption
Figure 4: Ablation study on the number of moment queries $\mathcal{K}$.
{adjustwidth}
-0.6cm0cm Refer to caption
Figure 5: Correlation between scores and ground-truth IoUs on QVHighlights val split.
[…]
Figure 6: Visualization of all moment span predictions for all the videos on QVHighlights val split, for all the 20 moment queries in the region-guided decoder.
[…]
Figure 7: Visualization of prediction comparisons on QVHighlights.
4.3 Ablation Study
To investigate the impact corresponding to key components of the proposed method, we conduct ablation studies on the validation set of QVHighlights.
[…]
Distribution of anchor initialization.  We adopt another simple uniform sampling strategy to replace the k-means algorithm. Specifically, we first generate a uniform grid on the normalized $m_{c}\times m_{\sigma}$ area, and uniformly sample 5 $\times$ 5 = 25 anchor pairs in a practical temporal region. As shown in Tab. 7, the performance drops significantly when replacing the k-means algorithm with the uniform sampling for explicit anchor initialization. It verifies that our k-means algorithm can provide optimal explicit regional priors for the decoding process.
IoU loss type.  As shown in Tab. 7, we employ different IoU loss to supervise the IoU score. All loss types can significantly improve performance, among which L2 loss achieves optimal performance.
Various scoring methods.  We utilize the product between classification confidence and IoU score as the final ranking criterion. Tab. 7 compares the product fusion with other methods, where IoU superv. means only using confidence score with IoU loss as supervision. All strategies have significant performance improvements, among which the product strategy achieves the best performance.
[…]
Correlation between scores and IoUs.  To compare IoU-aware scoring and classification confidence scoring, we draw scatter plots of the correlation between scores and ground-truth IoUs on the QVHighlights validation set in Fig. 5. It can be observed that our IoU-aware score shows a stronger correlation with the ground-truth IoU, i.e., the slope of the fitted line increases from 0.49 to 0.67, improving the distinction of high-quality proposals.
@@ 4.4 Visualization and Qualitative Result
As shown in Fig. 6, we visualize moment span predictions for all the 1550 QVHighlights val videos, for all the 20 moment queries in the region-guided decoder. Compared with previous methods in Fig. 1, RGTR introduces explicit regional guidance through anchor pairs, effectively eliminating numerous overlapped predictions and enhancing diversity.
In Fig. 7, we illustrate a qualitative example on QVHighlights, where the sentence corresponds to multiple moment spans. Since our method emphasizes enhancing prediction diversity, RGTR generates more accurate moment predictions than other methods, especially in the case of requiring simultaneous attention to different center coordinates and durations.
@@ 5 Conclusion
In this paper, we propose a Region-Guided TRansformer (RGTR) framework to address the limitations of DETR structure in the TSG task. To improve the diversity of queries, we design a region-guided decoder, which adopts a set of anchor pairs as moment queries to introduce explicit regional guidance for the decoding process. Each anchor pair takes charge of moment prediction for a specific temporal region, which reduces the optimization difficulty and ensures the diversity of the final predictions. To distinguish high-quality proposals, we employ an IoU-aware scoring head that incorporates localization quality to enhance classification confidence estimation. Extensive experiments on three channeling benchmarks demonstrate the superiority of our proposed RGTR.

### v2

@@ 3.3 Region-Guided Decoder
M-DETR (Lei, Berg, and Bansal 2021) 52.89 33.02 54.82 29.40 30.73 53.94 34.84 - - 32.20
QD-DETR (Moon et al. 2023b) 62.40 44.98 62.52 39.88 39.86 62.68 46.66 62.23 41.82 41.22
UniVTG (Lin et al. 2023) 58.86 40.86 57.60 35.59 35.47 59.74 - - - 36.13
TR-DETR (Sun et al. 2024) 64.66 48.96 63.98 43.73 42.62 67.10 51.48 66.27 46.42 45.09
TaskWeave (Yang et al. 2024) - - - - - 64.26 50.06 65.39 46.47 45.38
UVCOM (Xiao et al. 2024) 63.55 47.47 63.37 42.67 43.18 65.10 51.81 - - 45.79
CG-DETR (Moon et al. 2023a) 65.43 48.38 64.51 42.77 42.86 67.35 52.06 65.57 45.73 44.93
$\text{LLMEPET}^{\dagger}$ (Jiang et al. 2024) 66.73 49.94 65.76 43.91 44.05 66.58 51.10 - - 46.24
@@
Table 1: Performance Comparison on QVHighlights test and val splits. $\dagger$ indicates LLM-based method.
Method TACoS Charades-STA
R@0.3 R@0.5 R@0.7 mIoU R@0.3 R@0.5 R@0.7 mIoU
M-DETR (Lei, Berg, and Bansal 2021) 37.97 24.67 11.97 25.49 65.83 52.07 30.59 45.54
MomentDiff (Li et al. 2024) 44.78 33.68 - - - 55.57 32.42 -
UniVTG (Lin et al. 2023) 51.44 34.97 17.35 33.60 70.81 58.01 35.65 50.10
CG-DETR (Moon et al. 2023a) 52.23 39.61 22.23 36.48 70.43 58.44 36.34 50.13
$\text{LLMEPET}^{\dagger}$ (Jiang et al. 2024) 52.73 - 22.78 36.55 70.91 - 36.49 50.25
RGTR (Ours) 53.04 40.31 24.32 37.44 72.04 57.93 35.16 50.32
Table 2: Performances Comparison on TACoS and Charades-STA. $\dagger$ indicates LLM-based method.
Region-Guided Attention Module.  Similar to the general decoder, our region-guided decoder also includes two parts: self-attention module and cross-attention module. However, we employ different anchors in two modules for varying roles, as shown in Fig. 3. In the self-attention module, static anchors are utilized to focus content embeddings on preset representative temporal regions and share information across different regions. Specifically, we utilize static anchors as the positional embedding of self-attention module, such that the updated content embedding $C_{s}^{j}$ is as follows:
$C_{s}^{j}=\text{MultiHeadAttn}(C^{j-1}+P^{0}_{s},C^{j-1}+P^{0}_{s},C^{j-1}),$ (6)
where $C^{j-1}\in\mathbb{R}^{\mathcal{K}\times D}$ is the content embedding from ($j-1$)-th decoder layer, and $C^{0}$ is initialized to zeros. In the cross-attention module, we employ dynamic anchors as query positional embedding to aggregate region-specific features from fusion embedding $F$ with the assistance of $C_{s}^{j}$. Therefore, the content embedding is updated as:
$C^{j}=\text{MultiHeadAttn}([C^{j}_{s},P_{d}^{j}],\;[F,\text{PE}(F)],\;F),$ (7)
where [$\cdot$,$\cdot$] means concatenation function. By adopting anchor pairs with regional guidance, the decoder reduces the optimization difficulty and eliminates overlapped proposals.
3.4 IoU-Aware Scoring Head
The region-guided decoder improves the quality of proposals by reducing overlapped and redundant proposals, while high-quality proposals demand not only fewer duplications but also accurate boundaries. In the previous DETR-based methods (Jang et al. 2023; Sun et al. 2024), classification confidence (foreground or background) is adopted to rank all proposals. However, a single binary classification score may inadequately assess proposal quality by overlooking temporal boundary accuracy. To distinguish high-quality proposals, we introduce an IoU-aware scoring head, which considers both localization quality and classification confidence.
Specifically, the output of the decoder is fed to an FFN and a linear layer to predict the moment span and the confidence score $p_{c}$. Additionally, we add a linear layer to predict the expected IoU $p_{\text{IoU}}$. Instead of scoring proposals by classification confidence alone, we score them by a joint combination of confidence and IoU score, i.e., the product between $p_{c}$ and $p_{\text{IoU}}$. We supervise the IoU score with an L2 loss to the ground-truth IoU, denoted as $\hat{g}_{\text{IoU}}$,
This additional IoU score can explicitly incorporate localization quality to enhance classification confidence estimation, thereby generating high-quality proposals. Additionally, non maximum suppression (NMS) is applied during inference.
3.5 Training Objectives
The objective losses of RGTR include four parts: moment loss $\mathcal{L}_{\text{mom}}$, saliency loss $\mathcal{L}_{\text{sal}}$, alignment loss $\mathcal{L}_{\text{align}}$ and IoU loss $\mathcal{L}_{\text{IoU}}$. The overall objective is defined as:
where $\lambda_{\ast}$ are the balancing parameters. $\mathcal{L}_{\text{mom}}$ and $\mathcal{L}_{\text{sal}}$ are consistent with QD-DETR (Moon et al. 2023b).
4 Experiments
4.1 Datasets and Metrics
Datasets.  We evaluate the proposed method on three temporal sentence grounding benchmarks, including the QVHighlights (Lei, Berg, and Bansal 2021), Charades-STA (Gao et al. 2017), and TACoS (Regneri et al. 2013). QVHighlights spans various themes, Charades-STA comprises intricate daily human activities, and TACoS mainly showcases long-form videos focusing on culinary activities.
Metrics.  We adopt the Recall@1 (R1) under the IoU thresholds of 0.3, 0.5, and 0.7. Since QVHighlights contains multiple ground-truth moments per sentence, we also report the mean average precision (mAP) with IoU thresholds of 0.5, 0.75, and the average mAP over a set of IoU thresholds [0.5: 0.05: 0.95]. For Charades-STA and TACoS, we compute the mean IoU of top-1 predictions.
4.2 Implementation Details
Following previous methods (Moon et al. 2023b), we use SlowFast and CLIP to extract visual features and CLIP to extract textual features. We set the embedding dimension $D$ to 256. The number of anchor pairs $\mathcal{K}$ is set to 20 for QVHighlights, 10 for Charades-STA and TACoS. The NMS threshold is set to 0.8. The balancing parameters are set as: $\lambda_{\text{align}}=0.3$, $\lambda_{\text{iou}}=1$, and $\lambda_{\text{sal}}$ is set as 1 for QVHighlights, 4 for Charades-STA and TACoS. We train all models with batch size 32 for 200 epochs using the AdamW optimizer with weight decay 1e-4. The learning rate is set to 1e-4.
Method R0.5 R0.7 $\text{mAP}_{avg}$
Charades-STA-Len
2D-TAN (Zhang et al. 2020) 28.68 17.72 22.79
MMN (Wang et al. 2022b) 34.31 19.94 26.85
@@ 4.2 Implementation Details
MomentDiff (Li et al. 2024) 38.32 23.38 28.19
RGTR 61.17 40.23 44.30
Charades-STA-Mom
2D-TAN (Zhang et al. 2020) 20.44 10.84 17.23
MMN (Wang et al. 2022b) 27.20 14.12 19.18
@@
MomentDiff (Li et al. 2024) 33.59 15.71 21.37
RGTR 49.81 29.77 33.19
Table 3: Results on two out-of-distribution splits of Charades-STA. The VGG and Glove features are employed for all models. $\dagger$ indicates reproduced by official codebase.
Setting AEI RGAM IASH R0.5 R0.7 $\text{mAP}_{avg}$
(a) 65.35 48.97 43.12
(b) 64.65 50.58 44.82
(c) 66.19 49.61 44.03
(d) 65.55 51.29 45.36
(e) 66.13 51.68 46.51
(f) 67.68 52.90 46.95
Table 4: Ablation study on the components of RGTR on QVHighlights val split. It investigates the anchor explicit initialization (AEI), the region-guided attention module (RGAM), and the IoU-aware scoring head (IASH).
@@ 4.3 Performance Comparison
As shown in Tab. 1, we compare RGTR to previous methods on QVHighlights. For a fair comparison, we report numbers for both the test and validation splits. Our method achieves new state-of-the-art performance on almost all metrics. Specifically, RGTR outperforms the latest methods like LLMEPET (Jiang et al. 2024), achieving 67.12% at mAP@0.5 and 45.53% at $\text{mAP}_{avg}$ on the test split. On the validation split, RGTR also maintains its lead. The notable performance advantages of RGTR demonstrate the effectiveness of anchor pairs with explicit regional guidance.
Method Changes R0.5 R0.7 $\text{mAP}_{avg}$
Initialization Method random 66.19 49.61 44.03
uniform grid 67.10 50.97 44.93
k-means 67.68 52.90 46.95
Scoring Method IoU superv. 67.87 52.84 46.54
cls + IoU 67.23 52.39 46.92
cls × IoU 67.68 52.90 46.95
Table 5: Ablation study on initialization and scoring method.
Figure 4: Ablation study on number of moment queries $\mathcal{K}$.
Figure 5: Correlation between scores and ground-truth IoUs.
Tab. 2 presents comparisons on TACoS and Charades-STA. Our method achieves the best performance on TACoS. On Charades-STA, RGTR also maintains its competitiveness. However, we observe that while our results are notably superior on QVHighlights, the margin is slightly reduced on TACoS and Charades-STA. We attribute this to the biased distribution of the two datasets compared to QVHighlights, resulting in less query diversity learned by anchor pairs.
Figure 6: Visualization of moment predictions on QVHighlights val split, for all 20 dynamic anchors in region-guided decoder.
Figure 7: The qualitative result on QVHighlights.
4.4 Experiments on Out-of-Distribution Splits
To measure robustness, we also evaluate RGTR on two out-of-distribution splits (Li et al. 2024), Charades-STA-Len and Charades-STA-Mom, with distribution shifts of the length and moment location between training and test sets respectively. Since our anchor pairs are initialized by clustering centers on the training set, performance may degrade when the distribution changes significantly. However, as shown in Tab. 3, RGTR outperforms all previous methods under both out-of-distribution settings. Such surprising results indicate that the regional guidance introduced by anchor pairs works more by increasing the diversity between moment queries, rather than merely relying on the similarity between the training and test set distributions. Ablation experiments on other anchor initialization methods in Tab. 5 also confirm this point. Even with uniform grid points for initialization, which also serve as an initialization method to increase query diversity but are unrelated to the dataset distribution, the model performance improves significantly. Therefore, despite a distribution shift, the query diversity from regional guidance remains crucial for effective localization.
4.5 Ablation Study
@@
Anchor Initialization Method.  We adopt another two simple initialization methods to replace the k-means algorithm. “random” means utilizing random learnable queries as moment queries. “uniform grid” means generating a uniform grid on the normalized $m_{c}\times m_{\sigma}$ area, and uniformly sampling 5 $\times$ 5 = 25 points in a practical temporal region. As shown in Tab. 5, the performance of k-means initialization is significantly better than random initialization and uniform grid initialization. It verifies that k-means algorithm can provide optimal explicit regional priors for decoding process.
Scoring Method.  Tab. 5 compares the product fusion with other scoring methods, where IoU superv. means only using confidence score with IoU loss as supervision. All methods have significant performance improvements, among which the product method achieves the best performance.
@@
Correlation between Score and IoU.  To compare IoU-aware scoring and classification confidence scoring, we draw scatter plots of the correlation between scores and ground-truth IoUs on the QVHighlights validation set in Fig. 5. It can be observed that our IoU-aware score shows a stronger correlation with the ground-truth IoU, i.e., the slope of the fitted line increases from 0.49 to 0.67, improving the distinction of high-quality proposals.
@@ 4.6 Visualization and Qualitative Result
As shown in Fig. 6, we visualize all 20 dynamic anchors in the region-guided decoder on QVHighlights. Compared with previous methods in Fig. 1, RGTR introduces regional guidance through anchor pairs, effectively enhancing query diversity and eliminating numerous overlapped proposals.
In Fig. 7, we illustrate a qualitative example on QVHighlights, where the sentence corresponds to multiple moment spans. Since our method emphasizes enhancing query diversity, RGTR generates more accurate predictions than other methods, especially in the case of requiring simultaneous attention to different center coordinates and durations.
@@ 5 Conclusion
In this paper, we propose a Region-Guided TRansformer (RGTR) framework to address the limitations of DETR structure in TSG task. To eliminate overlapped proposals, we design a region-guided decoder, which adopts a set of anchor pairs as moment queries to introduce explicit regional guidance for decoding process. Each anchor pair takes charge of moment prediction for a specific temporal region, which reduces optimization difficulty and eliminates redundant proposals. To distinguish high-quality proposals, we employ an IoU-aware scoring head that incorporates localization quality to enhance classification confidence estimation. Experiments on three public datasets and two out-of-distribution splits demonstrate the superiority of RGTR.
Acknowledgments
This work was supported in part by National Science and Technology Major Project under Grant 2023ZD0121300, National Natural Science Foundation of China under Grants 62088102, 12326608 and 62106192, Natural Science Foundation of Shaanxi Province under Grant 2022JC-41, and Fundamental Research Funds for the Central Universities under Grant XTR042021005.